OpenStudy (austinl):

Find the general solution of the given differential equation, and use it to determine how solutions behave as \(t\rightarrow\infty\) \(y\prime + (1/t)y = 3\cos(2t)\) , \(t>0\)

OpenStudy (austinl):

@amistre64 Would you mind giving some guidance. I think I am probably making this more difficult than it should be.

OpenStudy (amistre64):

maybe in like an hour ill have more time to look at this

OpenStudy (austinl):

Ok, thanks very much!

OpenStudy (amistre64):

i know the homogenous is: c/t

OpenStudy (amistre64):

y = A(t)~t^(-1) y' = A'(t)~t^(-1) - A(t) t^(-2) ; let the A'(t) = 0 to account for the homogenous solution -A(t) t^(-2) = 3cos(2t) would most likely cover the particular ... is a thought.

OpenStudy (austinl):

Ok, the general solution of a problem is y=........ right? I believe we would have to do the \(\mu(t)\) thing right?

OpenStudy (amistre64):

im not adept at the u(t) thing .... never tried to work it like that

OpenStudy (amistre64):

the general solution would be: y = yh + yp in this case yh = C/t yp can be made with guesses, or by using variation of parameters or a wronskian if memory serves

OpenStudy (amistre64):

something like a guess of: yp = A cos(2t) + B sin(2t) is the usual run i believe

OpenStudy (austinl):

In my notes I think I just found something that may be helpful. \(\dfrac{dy}{dt}+ay=g(t)\) \(y=e^{-at}\int e^{at}g(t)dt+Ce^{-at}\) And now that I have it all written out, this particular formula only works if a is a constant. Darn it.

OpenStudy (austinl):

Ok ok. I think I found an example in my book that is somewhat similar to this problem \(2y\prime +ty=2\) Which can be written as, \(y\prime +\dfrac{t}{2}y=1\) Then using the \(\mu(t)\) method it says that, \(\mu(t)=e^{\int p(t)dt}=e^{\int \frac{t}{2}dt}=e^{\frac{t^2}{4}}\) Could I do something similar here?

OpenStudy (amistre64):

im just not that proficient or confident in the u(t) method to validate it

OpenStudy (austinl):

Ok, using words I can understand :P, would you mind walking me through your method and see if it is any better?

OpenStudy (amistre64):

in my head .... i something like this: y = A/t y/t = A/t^2 + y' = A'/t - A/t^2 ------------------ 3 cos(2t) = A'/t A' = 3t cos(2t) int A' = 3/4 (2t sin(2t) + cos(2t)) which is apprently off by that errant "t" bythe sin according to the wolf :)

OpenStudy (amistre64):

they say 2 sin(2t) + cos(2t)/t

OpenStudy (austinl):

I am sorry, I feel rather dumb right now, but I am having problems wrapping my head around this for some reason.

OpenStudy (austinl):

@Psymon Hate to randomly tag, but others have said that you may be able to help?

OpenStudy (abb0t):

This is a first order linear ODE. Start by obtaining your integrating factor.

OpenStudy (abb0t):

Once you get your integrating factor, distribute it across the ODE. You will see that you get product rule on the right hand side (guaranteed. unless you do your math wrong).

OpenStudy (austinl):

Ok, so I have: \(y\prime +(\dfrac{1}{t})y=3\cos(2t)\) Our integrating factor multiplied across: \((\mu(t))y\prime+(\dfrac{1}{t})(\mu(t))y=3(\mu(t))\cos(2t)\) As I stated earlier, \(\mu(t)=e^{\int p(t)dt}=e^{\int \frac{1}{t}dt}=e^{\ln|t|}\) Which I believe means that \(\mu(t)=t\) ?

OpenStudy (abb0t):

Yes, that is correct, \(\mu\)(t)= t Now, distribute \(\mu\)(t) across your ODE.

OpenStudy (abb0t):

You get: \(ty' +y = 3tcos(2t)\)

OpenStudy (abb0t):

Notice that your left hand side is nothing more than product rule! So you have: \([ty]' = 3tcos(2t)\)

OpenStudy (abb0t):

Integrate both sides with respect to \(t\) to solve for \(y\). The left side you have \(\int [yt]'dt\) = t(y(t)), now i WILL LET you integrate the right hand side. You will be using parts.

OpenStudy (austinl):

I am finding \(\int 3t\cos(2t)dt\) ?

OpenStudy (abb0t):

Yes, sir.

OpenStudy (austinl):

I see why I am integrating this side :P

OpenStudy (austinl):

\(=\dfrac{3}{4}(2t\sin(2t)+\cos(2t))+C\) Does this look correct?

OpenStudy (anonymous):

am I too late to this party?

OpenStudy (abb0t):

You can take over @oldrin.bataku I thought this was algebra 2 but it's not! So, bye.

OpenStudy (anonymous):

Find the general solution of the given differential equation, and use it to determine how solutions behave as t→∞ $$y'+\frac1ty=3\cos(2t)$$Recognize (hopefully by inspection) there is an obvious integrating factor \(\mu=t\) to use here:$$ty'+y=3\cos2t\\(ty)'=3\cos2t$$Integrating both sides we get:$$\int\frac{d}{dt}[ty]\,dt=\int 3\cos2t\,dt\\ty=3\int\cos2t\,dt=\frac32\sin(2t)+C\\y(t)=\frac3{2t}\sin2t+\frac{C}t$$

OpenStudy (austinl):

I believe you made an error in your second line of LaTeX, should it not be \(3t\cos(2t)\)?

OpenStudy (anonymous):

lol very good point

OpenStudy (austinl):

And may I ask, where does the "y" go on the left side?

OpenStudy (austinl):

Ok, so right now I am at \(ty=\dfrac{3}{4}(2t\sin(2t)+\cos(2t))\) Then do you just divide by t, and it is all good?

OpenStudy (austinl):

+C on end.... oops

OpenStudy (austinl):

+C on end.... oops

OpenStudy (abb0t):

C'mon gaiz...

OpenStudy (abb0t):

can't we all just be on the same tangent plane for once :-)

OpenStudy (anonymous):

i prefer jets

OpenStudy (austinl):

Ok, I think I have the answer. \(y(t)=\dfrac{3(2t\sin(2t)+\cos(2t))}{4t}+\dfrac{C}{t}\) Does this look even remotely correct?

OpenStudy (abb0t):

Ok mr.Fancy pants.

OpenStudy (abb0t):

and yes, that is correct.

OpenStudy (austinl):

WOOOOO!!!!!!! Then as \(t\rightarrow \infty\), does it approach zero? As there are t's in both denominators?

OpenStudy (abb0t):

Take the limit and find out.

OpenStudy (anonymous):

@austinL all of $$\sin2t,\cos2t,C$$are bounded while \(t\) is not so yes

OpenStudy (austinl):

When I take the limit, I arrive at \(\dfrac{-3}{2}~to~\dfrac{3}{2}\)

OpenStudy (anonymous):

that's because wolfram spit out lim sup/inf

OpenStudy (abb0t):

smh @wolfram

OpenStudy (austinl):

Zero then, go with my first answer?

OpenStudy (austinl):

I do believe I will work out another problem, close this and post the question and my solution and see if I am correct. Would you guys mind checking that in a few?